OpenStudy (anonymous):

How do you solve for 'x' from 73/8 + 6/2root(2) + x/root(2) + 6x + x^2 = 121/6 -8x + 4x^2 ?

OpenStudy (pfenn1):

Is this your equation?\[\frac{73}{8} + \frac{6}{2\sqrt{2}} + \frac{x}{\sqrt{2}} + 6x + x^2 = \frac{121}{6} -8x + 4x^2\]